OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be 3 interactions. If there are 3 children, there can be 9 interactions. If there are 4 children, there can be 18 interactions. How many interactions will there be for 9 children? A. 135 B. 84 C. 93 D. 108

OpenStudy (mathmath333):

the interactions are not shown in geometric progressions i think

OpenStudy (anonymous):

That's all there is to the question.

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

I think it's C

OpenStudy (phi):

why?

OpenStudy (phi):

one way to analyze this is find differences 1 0 3 2 3 3 6 3 9 3 9 4 18 the "second differences" are constant of 3 if it follows the same pattern the first differences are 0,3,6,9,12,15,18,21,24 and the nth entry is the sum of the first n terms. Example: the 3rd term would be 0+3+6 = 9 and the 4th term: 0+3+6+9= 18 0, 3, 9, 18, 30, 45, 63, 84, 108 1 2 3 4 5 6 7 8 9 a closed-form solution would be to notice that the nth term is 3 ( 0 + 1 + 2 + ... + n-1) = 3* (n-1)* n/2 for n=9: 3*8*9/2 =108